# Polynomial time sampling from log-smooth distributions in fixed dimension under semi-log-concavity of the forward diffusion with application to strongly dissipative distributions

### v2

@@ 2.2 Construction of our estimator
where the $y_{i}$ are independent Gaussians such that $y_{i}\sim\mathcal{N}(0,(1-e^{1-2t})I_{d})$ ; we shall refer to this estimator as self-normalized as common in the sampling literature [1]. Depending on the locations of the modes of $V$ we expect that this estimator behaves reasonably well in some regions of the space and poorly in others. Yet, the key property of self-normalized estimators is that they remain nearly bounded: in our case, it holds uniformly in $z,t$ that $\|\hat{s}_{t,n}(z)\|\leq\frac{\max_{i}\|y_{i}\|}{1-e^{-2t}}\sim\sqrt{d\log(n)}\,.$ This boundedness will allow us to derive a non-asymptotical control on the quadratic error that we present in the next section.
Remark 5
The estimator used in Ding et al., [18] is very similar to ours and simply differs by the fact that the authors discretize a Föllmer flow which corresponds to a reparametrized in time version of the standard OU process and where an arbitrary Gaussian may be targeted. The authors obtain strong guarantees yet under assumptions that exclude almost all distributions (see Appendix C). As a consequence, the authors avoid most of the technical difficulties we were confronted with.
3 Proof of the main result
In this section, we prove a non-asymptotic bound on the variance of the estimator presented above. First, we derive a general non-asymptotic bound that does not rely on any particular assumptions. Using the estimates provided in Mikulincer and Shenfeld, [36] and our additional semi-log-concavity assumption, this upper bound becomes fully explicit in the zeroth and second order moments of the ratio $\Phi_{t}=\frac{p_{t}^{2V}}{p_{t}^{V}}$. We then provide upper-bounds for these moments that again heavily leverage the estimates of Mikulincer and Shenfeld, [36]. Finally, we combine our results with those of Chen et al., 2023a [10] to state our main theorem.
3.1 Non asymptotical upper-bound of the variance
Before studying our ratio estimator defined in Eq. 4, we start by separately upper-bounding the variances of its numerator and denominator as a function of the forward process.
Proposition 6 (Variance of estimators)
Let $y_{1},\dots,y_{n}$ i.i.d. distributed as $\mathcal{N}(0,(1-e^{-2t})I_{d})$ and denote $\hat{N}(z)$ and $\hat{D}(z)$ the numerator and denominator of the estimator defined in Eq. 4, we have:
$\mathbb{E}[\|\hat{N}(z)-N(z)\|^{2}]$ $\leq\frac{p_{t}^{2V}Z_{2V}e^{-td}}{n}\left(\Delta\log(p_{t}^{2V})(z)-\frac{\Delta\log(\pi)(z)}{1-e^{-2t}}+\|\nabla\log(p_{t}^{2V})\|^{2}(z)\right)\,,$
$\mathbb{E}[\|\hat{D}(z)-D(z)\|^{2}]$ $\leq\frac{p_{t}^{2V}(z)Z_{2V}e^{-td}}{n}\,.$
The proof is deferred to Appendix A.2. We can now derive a non asymptotic quadratic error for our estimator.
Proposition 7 (Non-asymptotic bound of the quadratic error)
For all $z\in\mathbb{R}^{d}$, $n$ and $t>0$, it holds that
@@ 3.1 Non asymptotical upper-bound of the variance
The complete proof is left to Appendix A.3 yet we briefly sketch the main arguments. We split the expectation on the event $A$ where the empirical denominator $\hat{D}$ (resp. the empirical numerator $\hat{N}$) is not too small (resp. not too large) with respect to its expectation $D$ (resp. $\|N\|$) and on the complementary $\bar{A}$. Across $A$, we use a Taylor expansion of order $2$ to make the variance of the numerator and of the denominator appear and we then use the previous proposition. Conversely, the estimator remains almost bounded on $\bar{A}$. We use Tchebychev inequality to upper-bound $\mathbb{P}(\bar{A})$ and make the variances of the numerator and of the denominator appear again, which concludes the proof.
@@
Generic results on self-normalized estimators were derived in Agapiou et al., [1, Theorem 2.3]. Yet, it would have involved an additional $1/(p_{t}^{V})^{4}$ term which would have caused the integrated error $\int\mathbb{E}\left[\|\hat{s}_{t,n}(z)-\nabla\log(p_{t}^{V})(z)\|^{2}\right]\mathrm{d}p_{t}^{V}(z)$ to diverge.
Hence, we observe that the average quadratic error is related to the properties of the forward process Eq. 1 started at $\mu\propto e^{-V}$ and $\mu^{2}\propto e^{-2V}$ respectively. In particular, it involves the Laplacian of $\log(p_{t}^{2V})$ as well as the gradients of $\log(p_{t}^{V})$ and $\log(p_{t}^{2V})$. Hence, a first natural assumption is to bound the hessians of $\log(p_{t}^{V})$ and $\log(p_{t}^{2V})$ from above along the trajectory, a.k.a. semi-log-concavity along the forward.
Assumption 1 (Semi-log-concavity of the forward processes)
There exists $\beta\geq 0$ such that
$-\nabla^{2}\log(p_{t}^{V})\succeq-\beta I_{d}$ and $-\nabla^{2}\log(p_{t}^{2V})\succeq-\beta I_{d}$ uniformly over $\mathbb{R}^{d}$.
In order to control the gradient terms, we also need the converse bound across the whole trajectory. Fortunately, using the results of Mikulincer and Shenfeld, [36], smoothness at time $t=0$ suffices. We thus make the following additional assumption.
Assumption 2 (Smoothness)
The potential $V$ is $C^{2}(\mathbb{R}^{d})$ and has an $L$-Lipschitz gradient.
Proposition 9 (Mikulincer and Shenfeld, [36])
Under Assumption 2, it holds for all $z\in\mathbb{R}^{d}$ that
$\nabla^{2}\log(\pi)(z)-\nabla^{2}\log(p_{t}^{V})(z)\preceq\frac{(L-1)e^{-2t}}{(1-e^{-2t})(L-1)+1}I_{d}\,.$ (5)
The proof is left to Appendix A.4 and is mainly an identification of terms. As a result, the following lower bound holds: $\nabla^{2}\log(p_{t}^{V})\succeq-(L+1)I_{d}\,.$ In particular, under Assumptions 1-2, we recover that $\|\nabla\log(p_{t}^{iV})(z)\|\leq(iL+\beta+1)\|z\|+\|\nabla\log(p_{t}^{iV})(0)\|$ for $i\in\{1,2\}$. Hence, in order to control the average quadratic error integrated w.r.t. $p_{t}^{V}$, it suffices to control the zeroth and the second order moments of the ratio $\Phi_{t}(z)\coloneqq\frac{p_{t}^{2V}}{p_{t}^{V}}$. This will be the object of the next paragraph.
3.2 Estimates on the moments of the ratio
In this subsection, we provide our upper bounds on the moments of $\Phi_{t}$. By slight abuse of notations, we shall denote $m_{i}(\Phi_{t})=\int\|x\|^{i}\Phi_{t}(z)\mathrm{d}z$ the $i$-th moment of $\Phi_{t}$.
Lemma 10 (Bound on the moments of the ratio)
@@ 3.2 Estimates on the moments of the ratio
Assume that $\mu\propto e^{-V}$ is $L$-log-smooth and $C^{2}$ and has finite second moment. It holds that
$\begin{cases}&m_{0}(\Phi_{t})\leq\frac{(Z_{V})^{2}}{Z_{2V}}e^{td}(L(1-e^{-2t})+e^{-2t})^{d}\,,\\ &m_{2}(\Phi_{t})\leq\frac{2e^{t(d+2)}(Z_{V})^{2}}{Z_{2V}}(L(1-e^{-2t})+e^{-2t})^{d+2}\left[m_{2}(\mu)+d(L+1)+2d\log\left(\frac{L\mu(0)^{-2/d}}{2\pi}\right)\right]\,.\end{cases}$
The proof of Lemma 10 is provided in Appendix A.5 and relies on applying Grönwall’s lemma for the evolution of the moments. Crucially, note that this bound is only polynomial in $L$ when $d$ is fixed instead of exponential. This is a consequence of the bound of Mikulincer and Shenfeld, [36] that remains bounded with respect to $L$ whenever $t>0$. We have now all the ingredients to state our main result.
3.3 Assembling the pieces
We combine our non-asymptotic bound of Proposition 7 together with our previous estimates and Theorem 3 to recover a polynomial time bound in fixed dimension.
Theorem 11 (Main theorem)
@@ 3.3 Assembling the pieces
Under Assumptions 1-2 and the additional assumptions that there exists $\xi\geq 0$ such that $\|\nabla\log(p_{t}^{V})(0)\|,\|\nabla\log(p_{t}^{2V})(0)\|\leq\xi$ and that $\mu$ has finite second moment $m_{2}(\mu)$, if we run algorithm Eq. 3 for an appropriate (explicit) set of parameters and with the stochastic score estimators $\hat{s}_{n_{k},t_{k}}$ defined in Eq. 4, then, denoting $\hat{p}$ the stochastic distribution of the output $Y_{N}$, it holds that
$\begin{cases}&\mathbb{E}[\mathop{\mathrm{KL}}\nolimits(\mu,\hat{p})]\lesssim\epsilon\,,\\ &\sum_{i=1}^{N}n_{k}\lesssim d^{7}L^{d+2}\epsilon^{-2(d+3)}(m_{2}(\mu)+d)^{2(d+1)}(K_{1}+K_{2})\,,\\ &K_{1}=(L+\beta+1)^{2}(L+1)^{2}\bigg{[}m_{2}(\mu)+d(L+1)+2d\log\left(\frac{L\mu(0)^{-2/d}}{2\pi}\right)\bigg{]}\,,~{}~{}K_{2}=d\beta+\xi^{2}+1\,,\end{cases}$

### v3

@@ 4.2 Derivation of an estimator
where the $y_{i}$ are independent Gaussians such that $y_{i}\sim\mathcal{N}(0,(1-e^{1-2t})I_{d})$ ; we refer to this estimator as self-normalized as common in the sampling literature [Agapiou et al., 2017]. The key property of self-normalized estimators is that they remain nearly bounded: in our case, it holds uniformly in $z$ that
$\mathbb{E}[\|\hat{s}_{t,n}(z)]\|\leq\frac{\mathbb{E}[\max_{i}\|y_{i}\|]}{1-e^{-2t}}\sim\sqrt{\frac{d\log(n)}{1-e^{-2t}}}\,.$
This boundedness will allow us to derive a non-asymptotical control on the quadratic error that we present in the next section.
We now explain how this differs from previous work. While this estimator was already considered in [Saremi et al., 2024], the authors did not use it within the reverse diffusion pipeline and more importantly, we are the first work to derive quantitative guarantees for this estimator, which is one of our core contributions. Similarly, an estimator close to this one was considered in the context of Föllmer Flows [Jiao et al., 2021, Ruzayqat et al., 2023, Ding et al., 2023] in order to approximate the shift in the corresponding Schrödinger bridge. We show in Appendix B.1.1 that while this shift is itself close to the intermediate scores that we seek to approximate, the resulting self-normalized estimator is more degenerate. Namely, as discussed above, their guarantees are significantly weaker.
5 Sketch of proof of the main result
The proof is decomposed in three steps: (i) we derive a non-asymptotic bound on the quadratic error of the estimator presented in Eq. 5 (ii) we show that under Assumptions 1-2, the integrated error of this estimator (that appears in the bound of Theorem˜3) can be fully controlled by the zeroth and second order moments of the ratio $\Phi_{t}=p_{t}^{2V}/p_{t}^{V}$, where $p_{t}^{2V}$ (respectively $p_{t}^{V}$) is the density of the forward process defined in Eq. 1 initialized at $\mu^{2}$ (respectively $\mu$) (iii) we provide a quantitative bound on these moments as well as other relevant quantities and we conclude with Theorem 3.
Proposition 4 (Non-asymptotic bound on the quadratic error)
For all $z\in\mathbb{R}^{d}$, $n$ and $t>0$, denoting $p_{t}^{2V}$ (respectively $p_{t}^{V}$) the density of the forward process defined in Eq. 1 initialized at $\mu^{2}\propto e^{-2V}$ (respectively $\mu\propto e^{-V}$), $Z_{2V}$ (respectively $Z_{V}$) the normalizing constant of $\mu^{2}$ (respectively $\mu$) and $\pi$ the density of the standard Gaussian, it holds that
@@ 5 Sketch of proof of the main result
The complete proof is left to Appendix C yet we briefly sketch the main arguments. We split the expectation on the event $A$ where the empirical denominator $\hat{D}$ of Eq. 5 (respectively the empirical numerator $\hat{N}$) is not too small (respectively not too large) with respect to its expectation $D$ (respectively $\|N\|$) and on the complementary $\bar{A}$. Over $A$, we use a second-order Taylor expansion to make the variances of both the numerator and the denominator appear, and compute them explicitly. Conversely, the quadratic error of the estimator remains almost bounded on $\bar{A}$. We use Chebyshev’s inequality to upper-bound $\mathbb{P}(\bar{A})$ and make the variances of the numerator and of the denominator appear again, which concludes the proof.
@@
Generic bounds on self-normalized estimators were derived in Agapiou et al. [2017, Theorem 2.3]. Yet, we show in Appendix B.5 that if used in our context, they would involve at least an extra $1/(p_{t}^{V})$ factor which can cause the integrated error $\int\mathbb{E}\left[\|\hat{s}_{t,n}(z)-\nabla\log(p_{t}^{V})(z)\|^{2}\right]\mathrm{d}p_{t}^{V}(z)$ to diverge.
Then, we need to control the Laplacian of the forward processes as well as their gradient appearing in $\theta_{t}$ in the former proposition. As mentioned in the previous section, the intermediate scores can be re-written as
$\nabla\log(p_{t}^{V})(z)=\frac{z-e^{-t}\mathbb{E}_{q_{t,z}}[Y]}{1-e^{-2t}}\,$,
with $q_{t,z}(x)\propto e^{-V(x)}e^{-\frac{\|e^{-t}x-z\|^{2}}{2(1-e^{-2t})}}$. Now we note that if $\mu\propto e^{-V}$ is dissipative, so is $q_{t,z}$; in particular we can quantitatively bound its second order moment and a fortiori, upper-bound $\|\nabla\log(p_{t}^{V})(z)\|^{2}$.
Proposition 6 (Regularity bounds on the forward process)
Suppose that Assumption 2 holds. Then, for all $z\in\mathbb{R}^{d}$ and $t>0$, it holds that
$\begin{cases}&\|\nabla\log(p_{t}^{V})(z)\|^{2}\leq\frac{\|z\|^{2}}{(1-e^{-2t})^{2}}\biggl(2+\frac{e^{-2t}}{a(1-e^{-2t})}\biggr)+\frac{2e^{-2t}(2b+d)}{a(1-e^{-2t})^{2}}\,,\\ &\Delta\log(p_{t}^{V})(z)\leq\frac{e^{-2t}}{(1-e^{-2t})^{2}}\biggl(\frac{\|z\|^{2}}{2a(1-e^{-2t})}+\frac{2b+d}{a}\biggr)\,.\end{cases}$
The complete proof is left in Appendix D.2. This result implies that the $\theta_{t}$ term defined in Proposition 4 is at most of order of order $\theta_{t}(z)\sim(1+\|z\|^{2})$ w.r.t. $z$. In particular, the average integrated error $\mathbb{E}[\|\hat{s}_{t,n}(z)-\nabla\log(p_{t}^{V})\|_{L^{2}(p_{t}^{V})}^{2}]$ can be upper-bounded with respect to the zeroth and the second order moments of the ratio $\Phi_{t}=p_{t}^{2V}/p_{t}^{V}$. In the next proposition, we show that these moments are bounded under the semi-log-convexity assumption. By a slight abuse of notation, we shall denote $m_{i}(\Phi_{t})=\int\|x\|^{i}\Phi_{t}(z)\mathrm{d}z$ the $i$-th moment of $\Phi_{t}$.
Lemma 7 (Bounds on the moments of the ratio)
@@
Assume that $\mu\propto e^{-V}$ has finite second moment $m_{2}$ and that Assumption 1 holds. Then,
[…]
where $Z_{2V}$ (respectively $Z_{V}$) is the normalization constant of $e^{-2V}$ (respectively $e^{-V}$) and where $\pi$ in the right-hand-side refers to the constant $\pi\approx 3.14$.
The proof of this lemma is deferred to Appendix E. It relies on a key result in Mikulincer and Shenfeld [2023, Lemma 5] where it is shown that for $\beta$-semi-log convex distributions, the following bound holds:
$\nabla^{2}\log(\pi)-\nabla^{2}\log(p_{t}^{V})\preceq\frac{(\beta-1)e^{-2t}}{(1-e^{-2t})(\beta-1)+1}I_{d}\,.$
In particular, the right-hand-side remains bounded w.r.t. $\beta$ whenever $t>0$, which allows to avoid an exponential dependence in $\beta$ in our final bounds. It suffices now to control $\mathcal{I}(\mu,\pi),m_{2}$ and $\mu(0)$ in order to apply Theorem 3 and conclude.
Lemma 8
Assume that $\mu\propto e^{-V}$ is such that Assumption 1 and 2 hold: $\nabla^{2}V\preceq\beta I_{d}$ and $\langle\nabla V(x),x\rangle\geq a\|x\|^{2}-b$ for some $a>0,b,\beta\geq 0$. Then,
$\begin{cases}&m_{2}\leq(b+2d)/a\,,\\ &\mathcal{I}(\mu,\pi)\leq 2(b+2d)/a+2\beta d\,,\\ &\log(\mu(0)^{-2/d})\leq 4\beta b/(da)+2\pi+\log(2/a)\,.\end{cases}$
Here again, the $\pi$ in the right-hand-side refers to the constant $\pi\approx 3.14$.
The proof is left to Appendix D.1. We can now state our main result.
Theorem 9
@@
Under Assumptions 1 and 2, if we run the algorithm of Eq. 3 with $T=\log(1/{\epsilon})$, $N=1/{\epsilon}$, $t_{k}=kT/N$ and with the stochastic score estimators $\hat{s}_{n_{k},t_{k}}$ defined in Eq. 5 with $n_{k}=d\epsilon^{-(2d+3)}$, then, denoting $\hat{p}$ the stochastic distribution of the output $Y_{N}$, it holds that
$\mathbb{E}[\mathop{\mathrm{KL}}\nolimits(\mu,\hat{p})]\lesssim\epsilon\beta^{d+3}(b+d)/a^{2}\,$,